The question: The product of two negative numbers is 90. If x is one of the numbers, write their sum S as a function of x.
I'm not quite sure how it would be solved. The part that says "...their sum S as a function of x" is what confused me. Anyone who is willing to help is welcome to!

Answer by checkley77(12844)

90=x*y y is the other negative number.
x=90/y
S=x+y
S=x+90/y
x=S-90/y ans.